OpenStudy (anonymous):

Exponential equation 216^x-1=6

OpenStudy (anonymous):

what do you have to do solve or change to logarithm ?

OpenStudy (anonymous):

216 =6.6.6 so 6^3 then 6^3(x-1)=6, then 3x-3=1 ,x=4/3

OpenStudy (anonymous):

Thanks that was the answer i got when i worked it out the first time

OpenStudy (anonymous):

youre welcome